[Brahms-l] Shift report 20040319 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 19 2004 - 08:11:43 EST
------------------------------------------------------------
Date: Fri Mar 19 08:11:42 2004, Shift: 00:00-08:00
Supervisor: Michael Murray, with: Chaitanya Kalavagunta
Log book#15 pages 91 - 91

Shift Summary:
------------------------------------------------------------
1) Beam for about half shift 
2) Finished 8 deg A full field, changed magnet polarity but unfortunately left d2 off for runs 11135, 11136, 
Run 1137 is fine. 
TPC and normal pedestal runs

3) Network hung several times. We could not get supermon to begin taking data.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 19 08:12:04 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 19 2004 - 08:12:25 EST